OpenStudy (babynini):

Please check my work:) Suppose a particle moves back and forth along a straight line with velocity v(t), measured in feet per second and acceleration a(t).

OpenStudy (babynini):

a) what is the meaning of \[\int\limits_{60}^{120}v(t)dt\] b) what is the meaning of \[\int\limits_{60}^{120}\left| v(t) \right|dt\] c) what is the meaning of \[\int\limits_{60}^{120}a(t)dt\]

OpenStudy (babynini):

a) net change (how can I say this better..?) b) total distance traveled/total change in the interval 60 and 120 c) I'm not sure xD

OpenStudy (babynini):

@zepdrix :)

zepdrix (zepdrix):

Hmm I like part b :d

zepdrix (zepdrix):

for part a, maybe we use the word `displacement`. The integral tells us far away the particle is from it's position at t=60 to t=120. Somethinggggg like that... thinking...

OpenStudy (babynini):

hm hm "displacement of the particle" then?

zepdrix (zepdrix):

The integral in A tells us how far the particle has been displaced from its position at t=60 to its position at t=120. The integral in B tells us the total distance the particle traveled from its position at t=60 to its position at t=120.

zepdrix (zepdrix):

Ya, whatever wording makes sense to you :)

OpenStudy (babynini):

hehe your wording is much more eloquent.

zepdrix (zepdrix):

\[\large\rm \int\limits_{60}^{120}a(t)dt\quad=v(120)-v(60)\]So we still trying to make sense of this last part? :)

OpenStudy (babynini):

mhm.

zepdrix (zepdrix):

Hmm I guess we can't use the word displacement this time ;[ So maybe we go back to your original idea: Net change in velocity from t=60 to t=120.

OpenStudy (babynini):

What does that mean exactly?

OpenStudy (babynini):

I don't think a(t) = net change? o.o

zepdrix (zepdrix):

Let's say at t=60, the particle was moving 50mph. Then I guess at maybe t=70, it slows down to 30mph. And then how bout at t=90, it speeds up to 80mph. And then at t=120, it slows down to a velocity of 75mph. So the net change in velocity from t=60 to t=120 is the change from the start velocity to the end velocity: 75 - 50 = 25 From 50 to 30 is -20 from 30 to 80 is +50 from 80 to 75 is -5 -20 + 50 - 5 = 25

zepdrix (zepdrix):

That was a stupid explanation ^ But read it anyway :P See if it helps.

OpenStudy (babynini):

hrrm I am beginning to seee xD

zepdrix (zepdrix):

We can agree on this, yes?\[\large\rm \int\limits a(t)dt=v(t)\]So then if we give this integral limits we have,\[\large\rm \int\limits_{60}^{120}a(t)dt=v(120)-v(60)\]

zepdrix (zepdrix):

Yes :) but it should follow in the opposite direction, ya?\[\large\rm if\quad a(t)=v'(t),\qquad then\quad \int\limits a(t)=\int\limits v'(t)=v(t)\]

OpenStudy (babynini):

ou, yes :>

OpenStudy (babynini):

forgot about the integral being in front of it..haha..ha.

OpenStudy (babynini):

Yes, I agree!